# A Single Transformer for Scalable Vision-Language Modeling

### v1

@@ Abstract
We present SOLO, a single transformer for Scalable visiOn-Language mOdeling. Current large vision-language models (LVLMs) such as LLaVA mostly employ heterogeneous architectures that connect pre-trained visual encoders with large language models (LLMs) to facilitate visual recognition and complex reasoning. Although achieving remarkable performance with relatively lightweight training, we identify four primary scalability limitations: (1) The visual capacity is constrained by pre-trained visual encoders, which are typically an order of magnitude smaller than LLMs. (2) The heterogeneous architecture complicates the use of established hardware and software infrastructure. (3) Study of scaling laws on such architecture must consider three separate components — visual encoder, connector, and LLMs, which complicates the analysis. (4) The use of existing visual encoders typically requires following a pre-defined specification of image inputs pre-processing, for example, by reshaping inputs to fixed-resolution square images, which presents difficulties in processing and training on high-resolution images or those with unusual aspect ratio. A unified single Transformer architecture, like SOLO, effectively addresses these scalability concerns in LVLMs; however, its limited adoption in the modern context likely stems from the absence of reliable training recipes that balance both modalities and ensure stable training for billion-scale models. In this paper, we introduce the first open-source training recipe for developing SOLO, an open-source 7B LVLM with the single Transformer architecture using moderate academic resources (8 x A100 80GB GPUs). The training recipe involves initializing from LLMs, sequential pre-training on ImageNet and web-scale data, and instruction fine-tuning on our curated high-quality datasets. On extensive evaluation, SOLO demonstrates performance comparable to LLaVA-v1.5-7B, particularly excelling in visual mathematical reasoning [^1] The code is made public at https://github.com/Yangyi-Chen/SOLO..
@@ 1 Introduction
Large vision-language models (LVLMs) demonstrate remarkable performance on downstream tasks (Li et al., 2023c; Zhu et al., 2023; Liu et al., 2023c; Chen et al., 2023d; Kim & Ji, 2024). They can effectively extract visual information and follow human instructions to generate insightful responses (Li et al., 2023b). Two established approaches for vision-language modeling include: (1) Connecting pre-trained visual encoders (Dosovitskiy et al., 2021b; Radford et al., 2021) and large language models (LLMs) (Touvron et al., 2023; Jiang et al., 2023) via a learned projection module that maps the visual embeddings to the embedding space of LLMs (Dai et al., 2023; Gao et al., 2023; Liu et al., 2023c). (2) Leveraging a pre-trained visual encoder to extract features and aligning feature embeddings with a pre-defined codebook (Esser et al., 2021) to convert each image into a sequence of discrete visual tokens, thus enabling LVLMs to process both images and language tokens (Wang et al., 2022b; Peng et al., 2022; Anil et al., 2023; Team, 2024; Diao et al., 2023).
However, despite their effectiveness, these approaches have limitations that make them hard to scale. We consider an architecture scalable when it can demonstrate sustained performance improvement through scaling (e.g., more computational resources and/or data) without hitting any inherent bottleneck in the machine learning system. The bottleneck in prevalent LVLMs is evident across four dimensions (§2.1), primarily due to their reliance on a pre-trained visual encoder:
@@
    Challenges in large-scale training and deployment: The heterogeneous architecture of LVLMs with vision encoders poses challenges in adapting existing training and inference frameworks and utilizing existing hardware, which is mostly tailored and optimized for single Transformer architectures.
@@
    Limited image preprocessing flexibility complicates data scaling: Most vision encoders pre-define a specification on how image inputs should be pre-processed. For example, the widely used visual backbones, such as CLIP-ViT-336 (Radford et al., 2021), require a square image input with a resolution of $336\times 336$, making it challenging to scale LVLMs with such vision encoders by effectively training on more data with high-resolution images and those with unusual aspect ratios.
@@
While still lags behind recent state-of-the-art LVLMs on evaluation benchmarks, SOLO exhibits performance on par with LLaVA-v1.5-7B (§4) and the variant LLaVA-7B6), which is created following our training recipe in the controlled setting. In particular, SOLO distinguishes itself in the domain of visual mathematical reasoning. Furthermore, through comprehensive ablation studies, we validate the design choices of our training recipe. Our empirical results confirm that the sequential pre-training on ImageNet and web-scale datasets and instruction fine-tuning on our carefully curated data mixture are both essential for the training of such single Transformer LVLMs (§5). Interestingly, we find that after removing the first stage of pre-training on ImageNet, the LVLM will produce outputs of drastically different quality while exhibit similar image-conditioned language modeling loss (§5.1, Fig. 3).
@@ 2.1 Scalability Limitations in Existing LVLMs
Challenges in Large-Scale Training and Deployment
The heterogeneous architecture with multiple components complicates the implementation of machine learning systems for large-scale training and deployments. (1) Training challenge: At large training scale (i.e., multi-node clusters), it is necessary to distribute not only the Transformer-based LLMs but also the vision model and the MLP connector across multiple devices, employing techniques such as tensor and pipeline parallelism. Thus, prevalent LVLMs cannot directly use existing industry-grade training frameworks optimized for the Transformer architecture (Shoeybi et al., 2019; Cano et al., 2023), thus necessitating the development of new tensor-sharding mechanisms. In addition, AI alignment typically employs algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), which necessitate simultaneously maintaining multiple models (e.g., reward and critic models) in GPU memory and cause difficulty in the algorithm implementations for heterogeneous architectures. (2) Deployment: The heterogeneous architecture complicates the deployment process due to similar model and tensor sharding challenges described above. Consequently, this hampers the large-scale services of existing LVLMs. Moreover, existing specialized AI chips (Techcrunch, ) and inference libraries, such as vLLM (Kwon et al., 2023) and MLC-LLM (team, 2023), are mostly optimized for Transformer architectures, presenting significant challenges in the deployment of these models on end devices.
@@
The complexity introduced by the multiple components of LVLMs is a significant barrier to understanding and improving these systems. Each component —the visual encoder, the connector, and the language models—operates with its own set of parameters and training strategies (Radford et al., 2021; 2019; Brown et al., 2020a), which can lead to a lack of cohesion in the overall model behavior. Scaling laws are crucial for guiding the development of large foundational models by forecasting the performance of a target model using data from several sampled models that are significantly smaller in sizes (Kaplan et al., 2020; Bahri et al., 2021). However, applying these approaches to existing LVLMs requires simultaneous consideration and scaling of various components, hereby increasing complexity.
Limited Image Pre-Processing Flexibility Complicates Data Scaling
The strict requirements for image pre-processing imposed by the specifications of visual encoders may significantly impact the effective training of LVLMs on more data, hence complicating data scaling. For instance, the requirement of a consistent input resolution can make it difficult to process images that are naturally high-resolution or have non-standard aspect ratios without compromising on the quality or representational fidelity of the input. Current mitigation strategies involve splitting the original image into multiple sub-images, independently extracting visual features from each sub-image using pre-trained visual encoders, and subsequently aggregating the representation embeddings (Xu et al., 2024a; Liu et al., 2024a; Dong et al., 2024). However, these approaches seem ad-hoc and can be suboptimal as the backbone visual model is not pre-trained to handle these inputs, potentially impact the effective handling of these high-resolution images.
@@ 3.1 Model Architecture
def get_resize_output_image_size(image_size): l1, l2 = image_size if l2 ¡= l1: short, long = l2, l1 else: short, long = l1, l2
@@
if l2 ¡= l1: return new_long, new_short else: return new_short, new_long

### v3

# SOLO: A Single Transformer for Scalable Vision-Language Modeling

@@ Abstract
We present SOLO, a single transformer for Scalable visiOn-Language mOdeling. Current large vision-language models (LVLMs) such as LLaVA mostly employ heterogeneous architectures that connect pre-trained visual encoders with large language models (LLMs) to facilitate visual recognition and complex reasoning. Although achieving remarkable performance with relatively lightweight training, we identify four primary scalability limitations: (1) The visual capacity is constrained by pre-trained visual encoders, which are typically an order of magnitude smaller than LLMs. (2) The heterogeneous architecture complicates the use of established hardware and software infrastructure. (3) Study of scaling laws on such architecture must consider three separate components — visual encoder, connector, and LLMs, which complicates the analysis. (4) The use of existing visual encoders typically requires following a pre-defined specification of image inputs pre-processing, for example, by reshaping inputs to fixed-resolution square images. This inflexibility can create bottlenecks and impede scalability. A unified single Transformer architecture, like SOLO, effectively addresses these scalability concerns in LVLMs; however, its limited adoption in the modern context likely stems from the absence of reliable training recipes that balance both modalities and ensure stable training for billion-scale models. In this paper, we introduce the first open-source training recipe for developing SOLO, an open-source 7B LVLM with the single Transformer architecture using moderate academic resources (8 x A100 80GB GPUs). The training recipe involves initializing from LLMs, sequential pre-training on ImageNet and web-scale data, and instruction fine-tuning on our curated high-quality datasets. On extensive evaluation, SOLO demonstrates performance comparable to LLaVA-v1.5-7B, particularly excelling in visual mathematical reasoning [^1] The code is made public at https://github.com/Yangyi-Chen/SOLO..
@@ 1 Introduction
Large vision-language models (LVLMs) demonstrate remarkable performance on downstream tasks (Li et al., 2023c; Zhu et al., 2023; Liu et al., 2023c; Chen et al., 2023c; Kim & Ji, 2024). They can effectively extract visual information (Wang et al., 2023b) and follow human instructions to generate insightful responses (Li et al., 2023b; Chen et al., 2024d). Two established approaches for vision-language modeling include: (1) Connecting pre-trained visual encoders (Dosovitskiy et al., 2021b; Radford et al., 2021) and large language models (LLMs) (Touvron et al., 2023; Jiang et al., 2023) via a learned projection module that maps the visual embeddings to the embedding space of LLMs (Dai et al., 2023; Gao et al., 2023; Liu et al., 2023c), or an intermediate symbolic layer Wang et al. (2024c). (2) Leveraging a pre-trained visual encoder to extract features and aligning feature embeddings with a pre-defined codebook (Esser et al., 2021) to convert each image into a sequence of discrete visual tokens, thus enabling LVLMs to process both images and language tokens (Wang et al., 2022b; Peng et al., 2022; Anil et al., 2023; Team, 2024; Diao et al., 2023).
However, despite their effectiveness, these approaches have limitations that make them hard to scale. We define an architecture as scalable when it exhibits consistent performance improvements as computational resources and training data increase, maintaining a positive scaling law relationship up to practical limits. This is in contrast to architectures that show diminishing returns or performance plateaus at larger scales due to fundamental architectural limitations such as information bottlenecks (i.e., structural limitations in information transmission) in visual perception. The scalability limitations in prevalent LVLMs is evident across four dimensions (§2.1), primarily due to their reliance on a pre-trained visual encoder:
@@
    Challenges in efficient training and deployment: The heterogeneous architecture of LVLMs with vision encoders complicates adaptation to standard frameworks and hardware optimized for unified Transformer architectures, resulting in reduced computational efficiency in terms of training and inference speed.
@@
    Limited image pre-processing flexibility: Most vision encoders pre-define a specification on how image inputs should be pre-processed. For example, the widely used visual backbones, such as CLIP-ViT-336 (Radford et al., 2021), require a square image input with a resolution of $336\times 336$. The inflexibility of image pre-processing can cause bottlenecks that hinder scalability.
@@
While still lags behind recent state-of-the-art LVLMs on evaluation benchmarks, SOLO exhibits performance on par with LLaVA-v1.5-7B (§4) and the variant LLaVA-7B6), which is created following our training recipe in the controlled setting. In particular, SOLO distinguishes itself in the domain of visual mathematical reasoning. Further scalability analysis reveals SOLO’s better scaling behaviors, inference speed advantages, easier scaling laws analysis, and the scalability and benefits of our flexible image preprocessing pipeline (§6.2). In addition, through comprehensive ablation studies, we validate the design choices of our training recipe. Our empirical results confirm that the sequential pre-training on ImageNet and web-scale datasets and instruction fine-tuning on our carefully curated data mixture are both essential for the training of such single Transformer LVLMs (§5). Interestingly, we find that after removing the first stage of pre-training on ImageNet, the LVLM will produce outputs of drastically different quality while exhibiting similar image-conditioned language modeling loss (§5.1, Fig. 3).
@@ 2.1 Scalability Limitations in Existing LVLMs
Challenges in Efficient Training and Deployment
The heterogeneous architecture with multiple components complicates the implementation of machine learning systems for efficient training and deployments in terms of speed. (1) Training challenge: At large training scale (i.e., multi-node clusters), it is necessary to distribute not only the Transformer-based LLMs but also the vision model and the MLP connector across multiple devices, employing techniques such as tensor and pipeline parallelism. Thus, prevalent LVLMs cannot directly use existing industry-grade training frameworks optimized for the Transformer architecture (Shoeybi et al., 2019; Cano et al., 2023), thus necessitating the development of new tensor-sharding mechanisms. In addition, AI alignment typically employs algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), which necessitate simultaneously maintaining multiple models (e.g., reward and critic models) in GPU memory and cause difficulty in the algorithm implementations for heterogeneous architectures. (2) Deployment: The heterogeneous architecture complicates the deployment process due to similar model and tensor sharding challenges described above. Consequently, this hampers the large-scale services of existing LVLMs. Moreover, existing specialized AI chips (Techcrunch, ) and inference libraries, such as vLLM (Kwon et al., 2023) and MLC-LLM (team, 2023), are mostly optimized for Transformer architectures, presenting significant challenges in the deployment of these models on end devices.
@@
The complexity introduced by the multiple components of LVLMs is a significant barrier to understanding and improving these systems. Each component —the visual encoder, the connector, and the language models—operates with its own parameters and training strategies (Radford et al., 2021; 2019; Brown et al., 2020a), which can lead to a lack of cohesion in the overall model behavior. Scaling laws are crucial for guiding the development of large foundational models by forecasting the performance of a target model using data from several sampled models that are significantly smaller in sizes (Kaplan et al., 2020; Bahri et al., 2021). However, applying these approaches to existing LVLMs requires simultaneous consideration and scaling of various components, hereby increasing complexity.
Limited Image Pre-Processing Flexibility
The strict requirements for image pre-processing imposed by the specifications of visual encoders may create bottlenecks that hinder the scalability. For instance, the requirement of a consistent input resolution can make it difficult to process images that are naturally high-resolution or have non-standard aspect ratios without compromising on the quality or representational fidelity of the input. Current mitigation strategies involve splitting the original image into multiple sub-images, independently extracting visual features from each sub-image using pre-trained visual encoders and subsequently aggregating the representation embeddings (Xu et al., 2024a; Liu et al., 2024a; Dong et al., 2024). However, these approaches seem ad-hoc and can be suboptimal as the visual backbone is not pre-trained to handle these inputs, potentially impacting the effective handling of these high-resolution images.
@@ 3.1 Model Architecture
def get_resize_output_image_size(image_size): l1, l2 = image_size if l2 <= l1: short, long = l2, l1 else: short, long = l1, l2
@@
if l2 <= l1: return new_long, new_short else: return new_short, new_long
@@ 3.2.1 Pre-Training
Pre-Training Infrastructure
We modify the standard Megatron-LLM (Cano et al., 2023) to support arbitrary image patch inputs. We use one node with 8 NVIDIA A100 80G GPU for pre-training. We use 2-way tensor parallelism (Shoeybi et al., 2019) and 4-way data parallelism for training. Following Shoeybi et al. (2019), we adopt distributed optimizer to shard optimizer states across different GPUs for memory efficiency. In our test on an 8xA100 server with identical image-caption pre-training, Megatron achieves 20K tokens per second, 67% higher than DeepSpeed, making it more suitable for our pre-training requirements.
Training Hyperparameter
3.2.2 Instruction Fine-Tuning
Dataset Curation
We meticulously select a diverse range of supervised datasets to perform instruction fine-tuning, aiming to enhance their performance across various domains of vision-language tasks. Our dataset selection strategy is based on the empirical analysis derived in Laurençon et al. (2024); Lin et al. (2024); Lu et al. (2024), and is mainly driven by the objective to cover a comprehensive range of data types, including language-only data, detailed image captions, scientific documents, tables, documents, charts, OCR and text-rich images, and general visual question-answering (VQA) tasks. In §A, we present datasets and their statistics in Tab. 7 and more details regarding data curation.
Implementation Details
We utilize DeepSpeed (Rasley et al., 2020), as implemented in Accelerate (Gugger et al., 2022), for instruction fine-tuning. The choice to use Accelerate over Megatron for fine-tuning is based on practical efficiency. Accelerate enables quick experimentation with data mixtures by modifying a few lines of code, while Megatron demands extensive preprocessing for each run, complicating ablation studies. Additionally, Megatron’s complex, multi-layered implementation hinders customization, presenting challenges in potential further extension, such as introducing advanced methods like RLHF for alignment. For hyperparameters, the global batch size is configured at 640, with a weight decay parameter of 0.1. We train for 1 epoch with a maximum learning rate of 1e-5, which follows a linear warm-up phase and transitions to a cosine decay schedule.
4 Comparison to Existing LVLMs
4.1 Model
Each LVLM may have multiple variants based on different LLM sizes and architectures. If possible, we opt for the variant equipped with a 7B Mistral LLM. For the remaining LVLMs, we select the variant whose configuration most closely aligns with our specifications (Mistral-7B-LLM). We directly present the evaluation results of existing LVLMs from the leaderboard (OpenCompass, ) when available, to ensure a fair comparison.
4.2 Benchmarks
We select a wide range of benchmarks, encompassing both general vision-language tasks and specific task-oriented datasets, for evaluation and analysis. For general vision-language capability evaluation, we choose MMStar (Chen et al., 2024b), MME (Fu et al., 2024), and SEED-Bench (Li et al., 2024a). Specifically, MMStar measures elite vision-indispensable capabilities, MME measures both the perception and cognition capabilities, and SEED-Bench covers 12 evaluation dimensions covering various aspects of LVLMs capabilities. For scientific document understanding, we choose AI2D (Kembhavi et al., 2016) and ScienceQA (Lu et al., 2022a). For visual mathematical reasoning, we choose MathVista (Lu et al., 2023). We adopt VLMEvalKit (Contributors, 2023; Duan et al., 2024) to perform the unified evaluation.
Level Model Visual Language MMStar MME SEED ScienceQA MathVista AI2D
OpenFlamingo v2 (C) ViT-L/14 MPT-7B 26.9 607.2 28.8 44.8 18.6 31.7
MiniGPT-4-v2 EVA-G Llama2-13B 21.3 968.4 29.4 54.7 23.1 30.5
VisualGLM EVA-CLIP ChatGLM-6B 5.9 738.1 47.0 56.1 21.9 41.2
InstructBLIP EVA-G Vicuna-7B 32.7 1391.4 44.5 54.1 24.4 40.6
mPLUG-OWL v2 (C) ViT-L/14 Llama2-7B 34.8 1786.4 64.5 69.5 25.4 55.7
XComposer EVA-G InternLM-7B 6.9 1874.2 66.1 89.8 29.8 56.9
Level-2 MiniCPM-V SigLIP-400M MiniCPM-2.4B 38.6 1650.2 65.6 77.0 30.6 56.3
Monkey ViT-BigHuge Qwen-7B 37.0 1759.9 64.3 72.1 33.5 62.5
LLaVA-Next (C) ViT-L/14 Mistral-7B 38.4 1821.2 72.4 73.0 34.6 69.0
MiniCPM-v2 SigLIP-400M MiniCPM-2.4B 39.1 1808.2 67.1 80.7 39.8 62.9
Ours SOLO Mistral-7B 35.5 1260.0 64.4 73.3 34.4 61.4
4.3 Results
The experimental results are shown in Tab. 2. We find that SOLO significantly outperforms Level-1 LVLMs and also performs comparably to Level-2 LVLMs, despite slightly underperforming Level-3 LVLMs. Furthermore, SOLO excels in task-oriented benchmarks, especially in areas requiring scientific knowledge and mathematical reasoning, due to its successful integration of image representation and complex reasoning within a single unified model. Overall, while SOLO does not yet meet the SoTA performance of the leading LVLMs (Level-3) within the prevalent multi-component LVLM framework, it marks a substantial progression in unified vision-language modeling. It is important to consider that SOLO is trained using limited academic resources, specifically 8 A100 GPUs. This is in sharp contrast with the resources used to produce SoTA LVLMs. For example, the technical report of DeepSeek-VL (Lu et al., 2024) mentions that DeepSeek-VL-7B is trained on a cluster of 64 nodes, each comprising 8 Nvidia A100 GPUs, totaling 512 GPUs—this represents a resource scale 64 times greater than that used for SOLO. SOLO serves as a pivotal model, showcasing the scientific value of training LVLMs with unified architectures. This establishes SOLO as a viable candidate for future developments aimed at closing the performance gap with SoTA LVLMs, with more flexibility and scalability by avoiding issues in prior architectures (§2.1).
5 Validating Key Ingredients in Our Recipe
Select Checkpoints for Comparison
Qualitative Comparison
We randomly select one example for qualitative analysis (in Fig. 4). Despite the equivalent loss of the selected checkpoints in Fig. 3, we find that without ImageNet pre-training (Stage-2 only), the model generates irrelevant and meaningless image captions, indicates a training divergence.
Quantitative Comparison
Discussion
We hypothesize that discrepancies between a model’s vision and language capabilities can lead to the observed behaviors. Specifically, when there is a significant imbalance—such as with the Mistral 7B model, which possesses advanced language abilities but lacks vision understanding—the model may reduce loss by replicating caption patterns, including redundant text tokens irrelevant to the visual content. For instance, in a caption like “This is a dog”, the essential element is “dog”. Focusing solely on minimizing language modeling loss without a robust initialized vision representation may lead the model to favor generic phrases like “This is a" over the more discriminative “dog”. This is because the former includes more tokens, disproportionately influencing the overall language modeling loss. Pre-training on ImageNet at Stage 1, which emphasizes predicting only the “dog” token, helps the model develop a solid visual representation, effectively narrowing the gap between vision and language capabilities and mitigating this issue. In addition, the results indicate that pre-training loss on vision-language data does not reliably indicate the performance of LVLMs. Detailed analysis are provided in §7.1, which also demonstrates that training loss on the instruction fine-tuning data mixture is similarly unreliable as an indicator.
5.2 Stage-2 Pre-Training on Web-Scale Data
Stage-2 Pre-Training Improves Performance on Top of Stage-1
Combining ImageNet and Captioning at Stage-2 Hurts Performance
5.3 Performance Boost via Instruction Fine-Tuning
5.4 Additional Validation Experiments
6 Further Analysis
Figure 7: The controlled analysis of Fuyu-8B and LLaVA-7B.
6.1 Controlled Analysis of Fuyu and LLaVA
6.2 Scalability Analysis
Superior Scaling Properties of SOLO
We demonstrate that existing LVLMs with heterogeneous architectures exhibit diminishing returns despite increases in high-quality instruction fine-tuning data while SOLO shows better scaling behaviors. We perform instruction fine-tuning on pre-trained LLaVA obtained in §6.1 and compare its scaling behaviors with SOLO by measuring the performance improvement per training token. For evaluation, we fine-tune both models for 50 steps as their initial checkpoints, as their pre-trained versions are not effective at following instructions. Given that both models are fine-tuned on the same dataset, we can directly compare their average improvement in benchmark performance per training token during training. In the implementation, we measure performance improvement every 500 steps, normalized by the number of training tokens encountered during those steps, and average this to calculate the final metric. Fig. 9 shows that SOLO outperforms mLLaVA on the “performance improvement per token” metric across all evaluation benchmarks. This suggests SOLO benefits more from high-quality instruction fine-tuning data and demonstrates better scalability, indicating that its performance could further be improved more compared to mLLaVA with more data.
Figure 8: We compare the scaling behaviors of SOLO and LLaVA by measuring the improvement on benchmark performance per token.
Figure 9: The performance of SOLO when trained and tested on different resolutions of images. R@X denotes the resolution of X.
SOLO Exhibits Training and Inference Speed Advantage
Table 3: Inference latency comparison.
Model Inference Latency (seconds)
LLaVA-Next 0.0641
LLaVA-Interleave 0.0691
Qwen2-VL 0.0588
SOLO 0.0235
We measure the training speed of SOLO compared to LVLMs with heterogeneous architectures by measuring throughput. Using the same 8xA100 server, we compare the number of tokens processed per second during SOLO training and the official LLaVA implementation (Liu et al., 2023a). Our SOLO implementation achieves a throughput of 20K tokens per second, while LLaVA reaches 10.5K tokens per second, highlighting SOLO’s significant advantage in training speed. We also measure the inference speed of SOLO compared to several LVLMs with heterogeneous architectures, including LLaVA-Next, LLaVA-Interleave, and Qwen-VL. The evaluation is conducted on a single A100 GPU using 10,000 COCO images and a consistent prompt: “Generate the Caption for the <image>”. All inference code is implemented using HuggingFace, and the inference latency is measured from the input being provided to the model and the output of the first token for fair comparison among all models since they may generate free-form responses in different lengths. The results, shown in Tab. 3, indicate that SOLO consistently outperforms other LVLMs with heterogeneous architectures by a significant margin, demonstrating its clear advantage in inference speed and large-scale development.
SOLO Facilitates Easier Scaling Laws Analysis
Table 4: Predictability of the performance gains comparison between SOLO and mLLAVA on various benchmarks.
$R^{2}$ MMStar MME SEEDBENCH ScienceQA MathVista AI2D
mLLAVA 40.32 92.52 92.89 89.25 62.73 83.29
SOLO 88.75 92.36 98.59 97.75 84.70 96.55
We conduct a simplified scaling laws experiment to show that the performance gains from scaling up data for SOLO is more predictable compared to LLaVA-Style LVLMs. Specifically, we follow Kaplan et al. (2020) to fit the analytical function that LVLMs trained with a limited dataset (instruction fine-tuning in our case):
$L(D)=\left(\frac{D_{c}}{D}\right)^{\alpha_{D}},$
where $D_{c}$ and $\alpha_{D}$ are constants to be estimated, $D$ is the number of training tokens, $L(D)$ is the benchmark performance in our case. We use data points from the instruction fine-tuning of SOLO and mLLaVA to fit this function and report the coefficient of determination $R^{2}$, which reflects the quality of the fit and is equivalent to the predictability of performance. The results below demonstrate that SOLO ’s performance is more predictable compared to mLLaVA, indicating that SOLO facilitates easier scaling laws analysis. We further provide more discussion about the better scalability of SOLO in §7.3.
SOLO Demonstrates Improved Performance when Scaling up Image Resolution
We train SOLO on different image resolutions during the instruction fine-tuning stage for 1,000 steps due to the compute limits (see Fig. 9). The image resolution during inference matches that used in the instruction fine-tuning stage. We find that SOLO continues to improve the performance with increasing image resolution, especially for the visual mathematical reasoning task. In addition, there is no significant difference in performance between the 1024-square resolution and the adapted resolution with 1024 as the maximum used in SOLO. This demonstrates the efficiency and scalability of our flexible image pre-processing pipeline.
Table 5: Comparison of performance on ScienceQA and MathVista under conditions of extreme image resolutions.
Resolution > 800 ScienceQA MathVista Aspect Ratio > 3 ScienceQA MathVista
LLaVA-Next 69.53 30.18 LLaVA-Next 43.28 25.00
mLLAVA 67.09 28.87 mLLAVA 34.45 21.88
SOLO 71.19 32.02 SOLO 50.42 28.13
SOLO Benefits from its Dynamic High-Resolution Capability
We evaluate LVLMs on high-resolution images and images with extreme aspect ratios that diverge from those found in natural images. For a controlled analysis, we select two benchmark datasets (ScienceQA, MathVista) where SOLO and LLaVA-Next achieve comparable performance (within 1 absolute point). In addition, we implement mLLaVA, which follows LLaVA’s modeling framework but is trained using SOLO ’s recipe for further comparison. We select subsets from the original benchmarks that include images with either a width or height exceeding 800 pixels. Also, we select those subsets with a ratio (width/height) greater than 3 or less than 1/3 for extreme aspect ratio analysis. We evaluate LVLMs’ performance on these subsets. Our results, as shown in Tab. 5, verify SOLO ’s advantages regarding the performance on high-resolution images and images with extreme aspect ratios. The poorer performance of mLLaVA and LLaVA-Next is likely due to their heuristic resizing rules, which constrain images to predefined resolution settings. In contrast, SOLO retains the original aspect ratios, which appears to be a more optimal approach.
7 Discussion
(a) The training curves of two variants.
(b) The downstream performance of two variants.
Figure 10: The training curves and downstream performance evaluation of two variants of SOLO. We find that they show significant performance differences although achieving a similar loss on the instruction fine-tuning data mixture.
7.1 (Pre-)training Loss on Vision-Language Data is Not a Reliable Indicator of the Actual Performance
Figure 12: The evaluation of language capability.
Table 6: Comparison between SOLO and LVLMs with heterogeneous architectures regarding the complexity of scaling laws analysis.
Comparison SOLO LVLMs with Heterogeneous Architecture
Architecture Homogeneous Heterogeneous (visual encoder, connector, LLM)
Scaling laws analysis Standard approach (Hoffmann et al., 2022) Requires extra “parameter allocation law” (an additional factor in the scaling law formulation)
Key challenge N/A (widely explored in Hoffmann et al. (2022)) Optimal parameter allocation and consistency issues
Impact on computation Moderate High, due to multiple configurations and additional experiments for parameter allocation law
Risk of error propagation Low Higher due to additional complexities
7.3 Scaling Laws Analysis with SOLO
We demonstrate how SOLO ’s unified architecture facilitates easier analysis by comparing its experimental plan for deriving scaling laws to those of heterogeneous-architecture LVLMs. To simplify the analysis process, we follow the standard experimental setup in Hoffmann et al. (2022) to proportionally scale up the model size and training tokens by a factor of 20.
For SOLO, we can follow the standard scaling laws approach by selecting sampling models of varying sizes and assigning the appropriate number of training tokens based on the predefined scaling factor (e.g., 20 times). After training each model on its designated tokens, the pre-training loss can thus be measured to obtain the (expended FLOPs, performance) data point. The obtained data points can be used to fit the power law curve that predicts the performance of the target model (e.g., a model with 70B parameters) given the FLOPs expected to expend.
In contrast, for LVLMs with heterogeneous architectures, such as LLaVA, scaling laws analysis introduces additional complexity. Two key issues arise: (1) Parameter allocation: How should the model’s parameters be distributed among the three components (visual encoder, connector, LLM)? (2) Consistency during scaling: Will this parameter allocation (e.g., a 0.05:0.01:0.94 ratio) remain fixed as model size, data size, and compute scale up? To address these, an additional “parameter allocation law” must be derived to predict the optimal distribution of parameters based on model size and training tokens. In implementation, for each size of the sampling model, multiple configurations regarding parameter allocation must be tested to determine the optimal pre-training loss and the corresponding best allocation, significantly increasing computational demands. Moreover, it’s challenging to figure out the best analytical form to map the model size and training tokens to the parameter allocation ratios. Consequently, the introduction of the parameter allocation law complicates analysis and increases the risk of error propagation.
8 Related Work
Model Architecture
Training Data
Evaluation Benchmarks
The progress of LVLMs is guided and measured by the continuous development of evaluation benchmarks (Ferraro et al., 2015; Kafle et al., 2019; Gan et al., 2022; Chen et al., 2024d). Initially, evaluation primarily concentrates on fundamental visual-language skills, such as image captioning (Lin et al., 2014b; Plummer et al., 2015), basic visual information recognition (Antol et al., 2015; Goyal et al., 2017), compositional visual understanding (Hudson & Manning, 2019), and knowledge reasoning based on visual information (Marino et al., 2019; Schwenk et al., 2022). Current benchmarks are advancing to encompass more intricate capabilities, requiring LVLMs to perform detailed visual analysis and complex reasoning (Uppal et al., 2022; Zhang et al., 2024a). These benchmarks range from general assessments across various domains and skills (Li et al., 2024a; Fu et al., 2024; Chen et al., 2024b; Yu et al., 2023) to specific tests targeting particular abilities, such as scientific document understanding (Kembhavi et al., 2016; Lu et al., 2022a), mathematical reasoning (Lu et al., 2023; Wang et al., 2024a), multi-discipline understanding and reasoning (Yue et al., 2023; Wu et al., 2024), hallucination (Li et al., 2023e), and OCR ability (Liu et al., 2023d). In this work, we select the advanced general and skill-specific benchmarks for evaluation.
9 Conclusion
This work revisits the simple vision-language modeling framework with a single Transformer. We argue that this approach effectively mitigates the scalability limitations inherent in prevailing models. With academic resources, we build SOLO, a 7B LVLM initialized from the Mistral LLM. We detail the training recipe and conduct extensive analysis and evaluation to validate the ingredients in our recipe. Experimental results show that SOLO demonstrates performance comparable to LLaVA-v1.5, supporting the continued investigation into this unified vision-language modeling approach for improved scalability.
Limitations and Broader Impact Statement
The investigation into large-scale vision-language modeling using a unified transformer architecture remains nascent, with our model not yet reaching optimal performance across diverse benchmarks. Continued advancements in the direction of unified LVLMs for scalable vision-language modeling are anticipated. However, although developing LVLMs with strong capabilities brings significant advancements in AI, it also poses potential negative impacts. One concern is the risk of misuse, where the model could be employed for malicious purposes, such as generating misleading content that could manipulate public opinion or deceive individuals. Additionally, the model may inadvertently exacerbate biases present in the training data, leading to unfair or discriminatory outcomes in decision-making processes.
Acknowledgement
We thank the reviewers and the action editor for their valuable suggestions and comments. This research is based upon work supported by U.S. DARPA ECOLE Program No. HR00112390060. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Appendix
Appendix A Details of Instruction-tuning Data Curation